Show that the 2^n is not convergent.


1
Expert's answer
2021-05-31T08:10:16-0400

Solution:

We know that a necessary and sufficient condition for the convergence of a sequence is that it is bounded and has a unique limit point.

"2^{n}=(1+1)^{n} \geq 1+n, \forall n \in \mathbb{N}"

If P>0 is any real number howsoever large, we have,

1+n>P, whenever n>P-1 .

Let m be a positive integer, such that, m>P-1.

"\Rightarrow" For any real number P>0, "\exists" a positive integer, m, such that "2^{n}>P, \forall n \geq m"

"\Rightarrow" The sequence "\left\{2^{n}\right\}" is not bounded.

"\Rightarrow" The sequence "\left\{2^{n}\right\}" diverges to "\infty" .

"\Rightarrow" The sequence "\left\{2^{n}\right\}" is not convergent.

Hence, proved.